# Trojan Cleansing with Neural Collapse

### v2

@@ 1 Introduction
Trojan attacks, or backdoor attacks, aim to maliciously corrupt a neural network’s training process such that the presence of a trojan trigger in any input will force the network to produce a target classification [10]. This type of attack is generally accomplished by means of data poisoning, where the adversary adds their chosen trigger to a small portion of a training dataset and labels the poisoned samples with their selected target class. Models trained on this dataset then learn both the correct classifications for clean data and the malicious classification for inputs containing the trigger. Fig. 1, provided in [27], illustrates an example of clean data point (a) along a poisoned one (b). The yellow trigger in this example is clearly visible, but [25] showed that poisoning may take on a subtle nature that is undetectable by human auditors so that that malicious behavior can be hidden until test time or model deployment.
Figure 1: An example of trojan attack on a traffic sign classifier. (a) model works normally on clean data without triggers, (b) attackers manipulate the prediction by adding a specific trigger. Image due to [27].
In this paper, for the first time, we make a concrete connection between training with poisoned data and a phenomenon recently studied in the deep learning theory literature, called Neural Collapse [24]. Neural Collapse describes the widely observed phenomenon wherein, over the course of training a neural network, highly symmetric structures emerge in both the network’s weights and the feature representations of training samples. In contrast, trojan attacks are inherently asymmetric: they require any input be classified to a single target class given only the presence of a small perturbation. We hypothesize and experimentally demonstrate that the injection of a trojan trigger into a neural network breaks the symmetries of Neural Collapse. We then use these results to develop a carefully targeted, lightweight method for mitigating trojan attacks that is both architecture and dataset agnostic.
@@ 1.1 Contributions
There is a large body of work on trojan detection and mitigation, detailed in Sec. 2.2. Some of these approaches first attempt to understand how trojan triggers alter the structure of neural networks: Neural Cleanse [29] suggests that triggers activate a certain subset of neurons, the topological analysis of [35] finds that triggers introduce an unusual number of skip connections between early and late layers in the network. But these analyses are generally unable to predict precisely where and how a trigger will manifest in a particular neural network and must rely on data-driven approaches to identify, for instance, the particular subset of neurons which comprise the action of the trigger.
@@
We pursue this question through the inherent asymmetry introduced by training on poisoned datasets and the natural impact this has on the Neural Collapse (NC) phenomenon. We discuss NC in depth in Sec. 2.3, but, briefly, NC is a feature of training over-parameterized models where, as shown in Fig. 2, the final feature representations of the training data and the weight matrix of the final layer of the network, both, converge to a simple and highly symmetric geometric structure. We hypothesize that this symmetric structure, however, directly contradicts the asymmetry introduced by data poisoning. In this work, we present extensive experimental evidence supporting this hypothesis and show that the process of embedding a trojan in a neural network significantly weakens the NC phenomenon in measurable and predictive ways.
@@
We leverage these insights to initiate a new class of trojan cleansing methods: algorithms which take a potentially trojaned network and by imposing NC-inspired symmetry remove the trojan trigger while maintaining the network’s performance on clean data. In other words, identifying the predictive impact the trojan embedding has on NC metrics allows us to arrive at an entirely new method for cleansing. Many existing cleansing methods, as detailed in Sec. 2.2, rely either on some form of model compression, which may harm the model’s generalization capabilities, or on reconstructing the trojan trigger, which relies on knowledge of the class of the triggers the adversary may have chosen from. In contrast, we are able to introduce a small and carefully targeted adjustment to a small subset of the weights of the network without any knowledge of the possible trojan triggers: mitigating the attack without deteriorating the generalization power of the model.
@@
We conduct extensive evaluations of our cleansing method, comparing with many other standard cleansing algorithms over different network architectures and types of trojan attacks. Our experiments show that our cleansed network effectively maintains the accuracy of the original network on unseen, un-triggered test data, but is no longer susceptible to the trojan trigger. We achieve comparable performance to other cleansing algorithms against standard data poisoning attacks on ResNet architectures and state-of-the-art performance against more sophisticated trojan attacks and large transformer architectures. Additional experiments show our algorithm is particularly robust to data imbalance and corruption. Moreover, our algorithm is very lightweight and easy to implement for any standard classification network architecture. This stands in contrast to many other common cleansing techniques and is particularly important since the trojan threat model assumes that the user must outsource their model training and so likely has limited ML resources and expertise.
@@ 2 Related Work
Figure 2: The figure illustrates NC as training progresses from left to right. Green spheres represent the vertices of an ETF, red points represent the linear classifiers, large blue/green points represent class means, and small blue/green points represent final layer features, with different shades indicating different classes. As training progresses, the features concentrate. Figure due to [24].
@@ 2.1 Trojan Attacks
The baseline data poisoning trojan attack was explored in [5] and [10]. The TrojanNN [19] attack, on a pre-trained model, carefully designs a trigger via network analysis and then reinforces it with a data poisoning step. Wanet [23] introduces warping-based triggers, while ReFool [20] introduces reflection triggers, both offering stealthier triggers than traditional trojan attacks. Instead of using the manually defined trigger, Lira [8] jointly learns a subtle trigger injection function while poisoning the model.
Our work in this paper focuses on the image domain, but trojans have also been proven effective in other domains: [6] and [7] implement trojan attacks on language data and [32] poisons audio data to trojan a model for speech recognition. Extending our analysis and cleansing mechanism to these domains is an interesting direction for future work.
@@ 2.2 Model Cleansing
Li et al. [17] and Xia et al. [31] use forms of distillation to preserve only the model’s performance on clean data while Wang et al. [29] identifies and prunes neurons likely associated with the trojan trigger. Other approaches rely on first reconstructing potential triggers and then attempt to remove them: [4] learns a generative model to create likely triggers and then corrects the network via adversarial training. Zeng et al. [33] proposes a minimax formulation and computationally efficient optimization scheme for trigger reconstruction. Other approaches include [15], which leverages sparse reconstruction of inputs and feature representations to remove potential triggers, and [37], which applies adversarial training to a new layer inserted into the network. [36] is a detection and mitigation method that shrinks the norms of backdoor-related neurons by using sharpness-aware minimization with fine-tuning. Wu et al. [30] introduced BackdoorBench, a platform which implements many of these algorithms and provides a leader board that guides our benchmark selection in Sec. 4.2.
We note that our work is related but fundamentally distinct from trojan detection which aims to determine whether a model may be trojaned, including those focusd on data analysis, as in Baracaldo et al. [2], Liu et al. [18], Shen et al. [26] as well as those work that directly analyze the model, such as Chen et al. [3], Ma et al. [21], Fields et al. [9].
@@ 2.3 Neural Collapse
Neural Collapse is a phenomenon, first reported by Papyan et al. [24] and thoroughly reviewed by [16], which has been observed across a wide variety datasets and model architectures. It describes a process observed in over-parameterized classification deep neural networks which are trained past zero training error, into a period referred to as the terminal phase of training (TPT). As depicted in Fig. 2, as the TPT progresses, the final feature representations of the training data and the weight matrix of the network’s last layer converge to a highly uniform, symmetric geometric structure known as a simplex equiangular tight frame (simplex ETF).
@@
While NC is associated with models trained past perfect training accuracy, this practice is known to have practical benefits as in the double descent phenomenon [22]. Further experiments in [24] demonstrate in particular that the TPT offers advantages including improved generalization performance, enhanced robustness, and increased interpretability. Various theoretical models, based on locally elastic dynamics [34] and unconstrained features [38, 37] have begun examining how NC arises and its connections to these other neural network characteristics.
@@
It is important to note that Neural Collapse is not a binary phenomenon that either does or does not occur. In practice the characteristics of collapse arise even before the TPT and gradually converge over the course of the TPT. Consequently, our discussion of NC will focus on its degree, quantified by four metrics which converge to zero as the qualitative characteristics of NC obtain. We will define these metrics in Sec. 3.2 after introducing notation.
3 Neural Collapse Disruption Hypothesis
[24] and [12] observed and proposed metrics for Neural Collapse across diverse architectures and datasets. In this section, we review these metrics, and provide evidence that trojan triggers weaken Neural Collapse phenomenon. In particular, we evaluate various NC metrics associated with a variety of trojaned models and show that every trojaned model has at least one weaker NC metric. Furthermore, benign models consistently achieve at least the same degree of collapse as their trojaned counterparts across all metrics.
3.1 Neural Collapse: Definitions
We consider a deep learning classification problem with $K$ classes, $d-$dimensional training samples $\bm{X}_{i}{=}\left\{x_{1},\ldots,x_{n_{i}}\right\}$ for each class $i$, and neural network $f{:}\ \mathbb{R}^{d}{\to}\mathbb{R}^{K}$. Our analysis will focus on the final feature representation of the network, so we will decompose the network as $f(\bm{x})=\bm{W}g(\bm{x})+\bm{b}$, with $\bm{W}\in\mathbb{R}^{K\times m}$ and $\bm{b}\in\mathbb{R}^{m}$ giving the final fully-connected layer’s weights and biases, and $g:\mathbb{R}^{d}\to\mathbb{R}^{m}$ calculating the network’s final $m-$dimensional feature representation.
@@ 3.1 Neural Collapse: Definitions
The trojan attacks we evaluate in our experiments are carried out via data poisoning. To implement such an attack, we first choose a trigger function $\kappa:\mathbb{R}^{d}\to\mathbb{R}^{d}$. In our experiments this function will be the overlay of a small, solid patch on the base image, as shown in Fig. 1, or the application of an Instagram filter to the base image, as shown in Fig. 7. We then choose a poisoning proportion $\delta\in[0,1)$ and a target class $k_{T}\in[K]$. For each $\bm{X}_{k}$ with $k\neq k_{T}$ and each sample $\bm{x}\in\bm{X}_{k}$, with probability $\delta$ we delete $\bm{x}$ from $\bm{X}_{k}$ and add $\kappa(\bm{x})$ to $\bm{X}_{k_{T}}$. We say a model is trojaned if $\delta>0$, and a model is benign if $\delta=0$.
@@
[…]
@@
3.2 Neural Collapse: Metrics and Trojan Impact
The observed phenomenon of Neural Collapse can be more precisely described by a set of four characteristics laid out in [24] and quantified by a standard set of metrics defined in [12] and used in other analyses of NC such as [11] and [14]. In this section we will describe these metrics and present results comparing their convergence over the course of training for both benign and trojaned models on CIFAR-10, CIFAR-100, and GTSRB datasets. Our results show that the application of a trojan trigger causes the NC metrics to converge more slowly and to a larger value than their benign counterparts. There is variation in this behavior across different datasets and different metrics, but in every case the trojaned model shows weaker collapse.
Figure 3: Variability Collapse Comparison of the convergence of trojaned and benign models on the NC1 metric [Eq. 1] for GTSRB on ResNet18.
Figure 4: Self-duality Comparison of the convergence of trojaned and benign models on the NC3 metric [Eq. 3] for CIFAR-10 on ResNet18.
Figure 5: Nearest feature classification Comparison of the convergence of trojaned and benign models on the NC4 metric [Eq. 4] for GTSRB on ResNet18.
$\Sigma_{W}$ $\coloneqq\frac{1}{K}\sum_{k=1}^{K}\frac{1}{|\bm{X}_{k}|}\sum_{\bm{x}\in\bm{X}_{k}}\bigg{(}(g(\bm{x})-\bm{\mu}_{k})(g(\bm{x})-\bm{\mu}_{k})^{T}\bigg{)}$
$\Sigma_{B}$ $\coloneqq\frac{1}{K}\sum_{k=1}^{K}\bm{m}_{k}\bm{m}_{k}^{T}.$
[…]
Experimental evaluation of NC1 for both benign and trojaned models are shown in Fig. 5 and Tab. 1. On CIFAR-10 we did not observe a significant difference between the benign and trojaned models for this metric, but convergence of NC1 was significantly weaker in our experiments on CIFAR-100 and GTSRB. Note that the red dashed line in all figures represents the point at which the models achieve zero training error, and so the beginning of the terminal phase of training.
NC2: Convergence to a simplex ETF Both $\bm{M}$ and $\bm{W}$ converge to a simplex ETF. In particular, this means that the norms of each of the rows equalize and that they become maximally pairwise separated from each other as measured by cosine similarity. Note that the maximum pairwise cosine similarity among a set of $K$ vectors is $(1-K)^{-1}$. These two characteristics can be quantified, respectively, by the two statistics
[…]
(a) Equinorm metric [Eq. 2] for the feature class means and the final classifier layer weight matrix.
(b) Equiangular metric [Eq. 2] for the feature class means and the final classifier layer weight matrix.
@@ 3.2 Neural Collapse: Metrics and Trojan Impact
Fig. 6 tracks Eq. 2, which characterizes the convergence of both the classifier weights, $\bm{W}$ and the centered class means, $\bm{M}$ to a simplex ETF. For the equinorm property, shown in Fig. 6(a), the benign model shows the most uniformity in norm across both feature means and classifier weights, and as the percentage of trojaned data increases, the equinorm property weakens. This trend is particularly profound in the classifier weights. We note that, for both class means and classifiers, this is caused by the fact that the norm of the target class is consistently smaller than those of other classes. For the equiangular property a similar, though smaller trend is observed.
NC3: Convergence to self-duality The rows of the final layer weight matrix W converge to a simplex ETF dual to that of the feature representations. Given NC2, this means that $\bm{W}$ and $\bm{M}$ converge, up to a rescaling by their respective Frobenius norms.
[…]
@@
Fig. 5 shows this metric also converges to a smaller value for the benign model than either trojaned model and the disruption to Neural Collapse increases with the proportion of training data poisoned.
NC4: Nearest feature mean classification At inference time, the neural network uses a nearest neighbor decision rule in the final feature space. That is, a new test point is classified to the class with the feature mean nearest to its own feature representation. A metric for this is then the misclassification rate for this decision rule.
[…]
$\operatorname*{arg\,min}_{c\in[K]}\left\lVert g(\bm{x})-\bm{\mu}_{c}\right\rVert_{2}\bigg{)}$
@@
Similar to NC1, on CIFAR-10 we did not observe a significant difference between the benign and trojaned models for NC4 as the nearest neighbor rule reaches perfect accuracy for each of them. But collapse is noticeably weakened in this metric in our experiments on CIFAR-100 and GTSRB, as shown in Fig. 5 and Tab. 1.

### v3

@@ Abstract
keywords:
Trojan Cleansing, Adversarial Defense Mechanisms, Model Robustness, Backdoor Attacks, Neural Collapse
@@ 1 Introduction
Figure 1: An example of trojan attack on a traffic sign classifier. (a) model works normally on clean data without triggers, (b) attackers manipulate the prediction by adding a specific trigger. Image due to [tang2020embarrassingly].
Trojan attacks, or backdoor attacks, aim to maliciously corrupt a neural network’s training process such that the presence of a trojan trigger in any input will force the network to produce a target classification [badnets]. This type of attack is generally accomplished by means of data poisoning, where the adversary adds their chosen trigger to a small portion of a training dataset and labels the poisoned samples with their selected target class. Models trained on this dataset then learn both the correct classifications for clean data and the malicious classification for inputs containing the trigger. Figure 1, provided in [tang2020embarrassingly], illustrates an example of clean data point (a) along a poisoned one (b). The yellow trigger in this example is clearly visible, but [hiddentriggerbackdoorattacks] showed that poisoning may take on a subtle nature that is undetectable by human auditors so that that malicious behavior can be hidden until test time or model deployment.
In this paper, for the first time, we make a concrete connection between training with poisoned data and a phenomenon recently studied in the deep learning theory literature, called Neural Collapse [Papyan2020neural]. Neural Collapse describes the widely observed phenomenon wherein, over the course of training a neural network, highly symmetric structures emerge in both the network’s weights and the feature representations of training samples. In contrast, trojan attacks are inherently asymmetric: they require any input be classified to a single target class given only the presence of a small perturbation. We hypothesize and experimentally demonstrate that the injection of a trojan trigger into a neural network breaks the symmetries of Neural Collapse. We then use these results to develop a carefully targeted, lightweight method for mitigating trojan attacks that is both architecture and dataset agnostic.
@@ 1.1 Contributions
We pursue the trojan cleansing problem through the inherent asymmetry introduced by trojan attacks and the natural impact this has on the Neural Collapse (NC) phenomenon. We discuss NC in depth in section 2.3, but, briefly, NC is a feature of training over-parameterized models where, as shown in Fig. 2, the final feature representations of the training data and the weight matrix of the final layer of the network, both, converge to a simple and highly symmetric geometric structure. We hypothesize that this symmetric structure, however, directly contradicts the asymmetry introduced by data poisoning.
In this work, the contribution is mainly in three parts:
Firstly, we present extensive experimental evidence supporting this hypothesis and show that the process of embedding a trojan in a neural network significantly weakens the NC phenomenon in measurable and predictive ways.
@@
Secondly, we leverage these insights to initiate a trojan cleansing method: an algorithm which takes a potentially trojaned network and by imposing NC-inspired symmetry removes the trojan trigger while maintaining the network’s performance on clean data. Many existing cleansing methods, as detailed in section 2.2, rely either on some form of model compression, which may harm the model’s generalization capabilities, or on reconstructing the trojan trigger, which relies on knowledge of the class of the triggers the adversary may have chosen from. In contrast, we are able to introduce a small and carefully targeted adjustment to a small subset of the weights of the network without any knowledge of the possible trojan triggers: mitigating the attack without deteriorating the generalization power of the model.
@@
Finally, we conduct extensive evaluations of our cleansing method, comparing with many other standard cleansing algorithms over different network architectures and types of trojan attacks. Our experiments show that our cleansed network effectively maintains the accuracy of the original network on unseen, un-triggered test data, but is no longer susceptible to the trojan trigger. We achieve comparable performance to other cleansing algorithms against standard data poisoning attacks on ResNet architectures and state-of-the-art performance against more sophisticated trojan attacks and large transformer architectures. Additional experiments show our algorithm is particularly robust to data imbalance and corruption. Moreover, our algorithm is very lightweight and easy to implement for any standard classification network architecture. This stands in contrast to many other common cleansing techniques and is particularly important since the trojan threat model assumes that the user must outsource their model training and so likely has limited ML resources and expertise.
@@ 2.1 Trojan Attacks
Early data poisoning trojan attacks were presented in [chen2017targeted] and [badnets]. The TrojanNN [liu2018trojann] attack, on a pre-trained model, customizes a trigger for the network and then reinforces it with a data poisoning step. Wanet [nguyen2021wanet] uses warping-based triggers and ReFool [liu2020reflection] uses reflection triggers, with both intended to produce subtler triggers than traditional trojan attacks. Instead of using a manually defined trigger, Lira [doan2021lira] simultaneously learns a trigger injection function while poisoning the model.
Our work in this paper focuses on the image domain, but trojans have also been proven effective in other domains: [badnl] and [lstmBasedTextClass] implement trojan attacks on language data and [audiotrojan] poisons audio data to trojan a model for speech recognition. Extending our analysis and cleansing mechanism to these domains is an interesting direction for future work.
@@ 2.2 Model Cleansing
[NeuralAttentionDistillation] and [kd2] use forms of distillation to preserve only the model’s performance on clean data while [neuralCleanse] identifies and prunes neurons likely associated with the trojan trigger. Other approaches rely on first reconstructing potential triggers and then attempt to remove them: [deepInspect] learns a generative model to create likely triggers and then corrects the network via adversarial training. [zeng2021adversarial] proposes a minimax formulation and computationally efficient optimization scheme for trigger reconstruction. Other approaches include [cleann], which leverages sparse reconstruction of inputs and feature representations to remove potential triggers, and [zhu2024neural], which applies adversarial training to a new layer inserted into the network. [zhu2023enhancing] is a detection and mitigation method that shrinks the norms of backdoor-related neurons by using sharpness-aware minimization with fine-tuning. [wu2022backdoorbench] introduced BackdoorBench, a platform which implements many of these algorithms and provides a leaderboard that guides our benchmark selection in section 4.2.
We note that our work is related but fundamentally distinct from trojan detection which aims to determine whether a model may be trojaned, including those focused on data analysis, as in [bacaraldoDetectingPoisoning, liuAnomolyDetection, shenPoisonDetection], as well as those works that directly analyze the model, such as [chenActivationDetection, maActivationDetection, fieldsDetection].
@@ 2.3 Neural Collapse
2.3.1 NC Overview
Figure 2: The figure illustrates NC as training progresses from left to right. Green spheres represent the vertices of an ETF, red points represent the linear classifiers, large blue/green points represent class means, and small blue/green points represent final layer features, with different shades indicating different classes. As training progresses, the features concentrate. Figure comes from [Papyan2020neural].
@@ 2.3.1 NC Overview
Neural Collapse is a phenomenon, first reported by [Papyan2020neural] and thoroughly reviewed by [kothapalli2022neural], which has been observed across a wide variety datasets and model architectures. It describes a process observed in over-parameterized classification deep neural networks which are trained past zero training error, into a period referred to as the terminal phase of training (TPT). As depicted in fig. 2, as the TPT progresses, the final feature representations of the training data and the weight matrix of the network’s last layer converge to a highly uniform, symmetric geometric structure known as a simplex equiangular tight frame (simplex ETF).
@@
While NC is associated with models trained past perfect training accuracy, this practice is known to have practical benefits as in the double descent phenomenon [nakkiran2021double]. Further experiments in [Papyan2020neural] demonstrate in particular that the TPT offers advantages including improved generalization performance, enhanced robustness, and increased interpretability. Various theoretical models, based on locally elastic dynamics [zhang2021imitating] and unconstrained features [zhu2021geometric, zhu2024neural] have begun examining how NC arises and its connections to these other neural network characteristics.
@@
It is important to note that Neural Collapse is not a binary phenomenon that either does or does not occur. In practice, the characteristics of collapse arise even before the TPT and gradually converge over the course of the TPT. Consequently, our discussion of NC will focus on its degree, quantified by four metrics which converge to zero as the qualitative characteristics of NC obtain. We will define these metrics in section 2.3.3 after introducing notation.
2.3.2 NC Definitions
[…]
@@ 2.3.2 NC Definitions
The trojan attacks we evaluate in our experiments are carried out via data poisoning. To implement such an attack, we first choose a trigger function $\kappa:\mathbb{R}^{d}\to\mathbb{R}^{d}$. In our experiments this function will be the overlay of a small, solid patch on the base image, as shown in fig. 1, or the application of an Instagram filter to the base image, as shown in section 4.3.2. We then choose a poisoning proportion $\delta\in[0,1)$ and a target class $k_{T}\in[K]$. For each $\bm{X}_{k}$ with $k\neq k_{T}$ and each sample $\bm{x}\in\bm{X}_{k}$, with probability $\delta$ we delete $\bm{x}$ from $\bm{X}_{k}$ and add $\kappa(\bm{x})$ to $\bm{X}_{k_{T}}$. We say a model is trojaned if $\delta>0$, and a model is benign if $\delta=0$.
@@
2.3.3 NC Metrics
The observed phenomenon of Neural Collapse can be more precisely described by a set of four characteristics laid out in [Papyan2020neural] and quantified by a standard set of metrics defined in [han2021neural] and used in other analyses of NC such as [guo2024cross] and [hong2024neural].
$\begin{split}&\Sigma_{W}\coloneqq\frac{1}{K}\sum_{k=1}^{K}\frac{1}{|\bm{X}_{k}|}\sum_{\bm{x}\in\bm{X}_{k}}\bigg((g(\bm{x})-\bm{\mu}_{k})(g(\bm{x})-\bm{\mu}_{k})^{T}\bigg),\quad\\ &\Sigma_{B}\coloneqq\frac{1}{K}\sum_{k=1}^{K}\bm{m}_{k}\bm{m}_{k}^{T}.\end{split}$
$\textbf{NC1}\coloneqq\frac{1}{K}\operatorname{tr}\{\Sigma_{W}\Sigma_{B}^{T}\}.$ (1)
NC2: Convergence to a simplex ETF Both $\bm{M}$ and $\bm{W}$ converge to a simplex ETF. In particular, this means that the norms of each of the rows equalize and that they become maximally pairwise separated from each other as measured by cosine similarity. Note that the maximum pairwise cosine similarity among a set of $K$ vectors is $(1-K)^{-1}$. These two characteristics can be quantified, respectively, by two statistics:
$\begin{split}&\textbf{NC2}_{\text{Norm}}(\textbf{V})\coloneqq\frac{\underset{k\in[K]}{\operatorname{std}}(\left\lVert\textbf{v}_{k}\right\rVert_{2})}{\underset{k\in[K]}{\operatorname{mean}}\left(\left\lVert\textbf{v}_{k}\right\rVert_{2}\right)},\\ &\textbf{NC2}_{\text{Angle}}(\textbf{V})\coloneqq\underset{i,j\in[K],i<j}{\operatorname{mean}}\left(\left|\cos{\left(\textbf{v}_{i},\textbf{v}_{j}\right)}+\frac{1}{K-1}\right|\right).\end{split}$ (2)
NC3: Convergence to self-duality The rows of the final layer weight matrix W converge to a simplex ETF dual to that of the feature representations. Given NC2, this means that $\bm{W}$ and $\bm{M}$ converge, up to a rescaling by their Frobenius norms.
$\textbf{NC3}\coloneqq\left\lVert\frac{\textbf{W}}{\left\lVert\textbf{W}\right\rVert_{F}}-\frac{\textbf{M}}{\left\lVert\textbf{M}\right\rVert_{F}}\right\rVert^{2}$ (3)
NC4: Nearest neighbor classification At inference time, the neural network uses a nearest neighbor decision rule in the final feature space. That is, a new test point is classified to the class with the feature mean nearest to its own feature representation. A metric for this is then the misclassification rate for this decision rule.
$\begin{split}\textbf{NC4}\coloneqq&\frac{1}{K}\sum^{K}_{k=1}\frac{1}{|\bm{X}_{k}|}\sum_{\bm{X}\in\bm{X}_{k}}\mathbb{I}\bigg(\operatorname*{arg\,max}_{c\in[K]}\textbf{w}_{c}^{T}g(\bm{x})+b_{c}\\ &\neq\operatorname*{arg\,min}_{c\in[K]}\left\lVert g(\bm{x})-\bm{\mu}_{c}\right\rVert_{2}\bigg)\end{split}$ (4)
3 Trojans Disrupt Neural Collapse
In this section, we present results comparing the NC metrics’ convergence over the course of training for both benign and trojaned models on CIFAR-10, CIFAR-100, and GTSRB datasets. Our results show that the application of a trojan trigger causes the NC metrics to converge more slowly and to a larger value than their benign counterparts. There is variation in this behavior across different datasets and different metrics, but in every case the trojaned model shows weaker collapse.
(a) Equinorm metric 2 for the feature class means and the final classifier layer weight matrix.
(b) Equiangular metric 2 for the feature class means and the final classifier layer weight matrix.
Figure 4: Variability Collapse Convergence of NC1 metric [1] for trojaned and benign models for ResNet18 on GTSRB.
Figure 5: Self-duality Convergence of the NC3 metric [3] for trojaned and benign models for ResNet18 on CIFAR-10.
Figure 6: NN Classification Convergence of the NC4 metric [4] for trojaned and benign models for ResNet18 on GTSRB.
@@ 3 Trojans Disrupt Neural Collapse
$\text{NC2}_{\text{Norm}}(\bm{M})$ 0.0325 0.0372 0.0453 0.1679 0.1666 0.1905 0.1775
$\text{NC2}_{\text{Norm}}(\bm{W})$ 0.0089 0.0169 0.0527 0.1796 0.2166 0.2584 0.2798
$\text{NC2}_{\text{Angle}}(\bm{M})$ 0.0511 0.0507 0.0513 0.0719 0.0862 0.0777 0.0952
$\text{NC2}_{\text{Angle}}(\bm{W})$ 0.0518 0.0513 0.0520 0.0625 0.0778 0.0651 0.0988
@@
Figure 3 tracks NC2, defined in Equation 2, which characterizes the convergence of both the final classifier weights and the training feature means to a simplex ETF. The red dashed line in the figures marks the point where the training error reaches zero and so the regime in which Neural Collapse is generally observed. For the equinorm property, shown in Figure 3a, the benign model shows the most uniformity in norm across both feature means and classifier weights, and as the percentage of trojaned data increases, the equinorm property weakens. This trend is particularly profound in the classifier weights. We note that, for both class means and classifiers, the source of this deviation is actually that the norm of the target class is consistently smaller than those of other classes. For the equiangular comparison, a similar, though smaller trend is observed. In Figure 3b, the benign model reaches the smallest equiangular metric for both feature class means and classifiers. This further underscores the weakening of Neural Collapse due to trojaning.
@@
NC3, defined in Equation 3, is shown in Figure 6. Here, again, the metric converges to a smaller value for the benign model than either trojaned model and the disruption to Neural Collapse appears to increase with the proportion of training data poisoned with the trojan trigger.
@@
But this pattern is not observed in every collapse metric for every dataset. On CIFAR-10 we did not observe a significant difference between the benign and trojaned models for NC1, characterizing the variability across feature representations, or NC4, characterizing how close the network’s decision rule is to a nearest neighbor rule in the final feature space, as they all go down to zero. However, collapse is noticeably weakened in these metrics in our experiments on CIFAR-100 and GTSRB, as shown in Figures 6 and 6 and Table 1.
Thus, while not every metric clearly differentiates trojan and benign models on every dataset, it is the case that benign models always show at least the same degree of convergence as any trojaned model on every collapse metric. And, importantly, on every dataset we observe at least one metric where trojaned models show a significantly weaker degree of collapse than their benign counterpart–suggesting strong tension between Neural Collapse and an effective trojan trigger.